Colleen and Kendra start out 106.4 miles apart. They begin riding their bicycles toward each other with Colleen riding five mph faster than Kendra. If they meet in four hours, how fast was each person riding?

Answer by lwsshak3(11628)

Colleen and Kendra start out 106.4 miles apart. They begin riding their bicycles toward each other with Colleen riding five mph faster than Kendra. If they meet in four hours, how fast was each person riding?
let x=speed at which Kendra is traveling
x+5=speed at which Coleen is traveling
They are traveling toward each other at the sum of their indv. speeds, x+x+5=2x+5
speed=distance/travel time
106.4/(2x+5)=4
8x+20=106.4
8x=86.4
x=10.8
x+5=15.8
=speed at which Kendra is traveling=10.8 mph
x+5=speed at which Coleen is traveling=15.8 mph